# Subtlety of oscillation indices of oscillatory integrals of real analytic functions

### v1

@@ Abstract.
For a locally defined real analytic function, we study the relation between the oscillation index of oscillatory integrals and the real log canonical threshold. The former is always negative, and its absolute value is greater than or equal to the latter. They coincide very often, but there are certain exceptional cases even in the Newton nondegenerate convenient homogeneous case where the number of variables is even and smaller than the degree.
@@ Introduction
Note that the normal crossing case is a very special case of a Newton nondegenerate case, see also [CKN 13]. It seems to be believed by some people that the equality holds in (5) at least in the Newton nondegenerate case with $\gamma(f)\,{<}\,1$.
@@
Note that if $f$ is homogeneous and $d$ is odd, then $\dim f^{-1}(0)\,{=}\,n{-}1$, see Remark  3.1 below. (Theorem  3 does not seem quite compatible with some well-known assertion in the literature, see for instance [CKN 13].)

### v2

@@ Abstract.
For a locally defined real analytic function, we study the relation between the oscillation index of oscillatory integrals and the real log canonical threshold. The former is always negative, and its absolute value is greater than or equal to the latter. They coincide very often, but there are certain exceptional cases even in the Newton nondegenerate convenient homogeneous case, for instance if the number of variables is even and smaller than the degree. This does not seem compatible with some standard formula in the literature, and there must be some error somewhere, although it does not seem easy to find it inside this paper.
@@ Introduction
Note that the normal crossing case is a very special case of a Newton nondegenerate case, see also [CKN 13]. It seems to be believed by some people that the equality holds in (5) at least in the Newton nondegenerate case with $\gamma(f)\,{<}\,1$, see for instance [AGV 88], [Va 76].
@@
Note that if $f$ is homogeneous and $d$ is odd, then $\dim f^{-1}(0)\,{=}\,n{-}1$, see Remark  3.1 below. Theorem  3 does not seem compatible with some well-known formula in the literature, see for instance [Va 76], [CKN 13]. So there must be some error somewhere, although it does not seem easy to find it inside this paper.